Exercise 6-28 Equivalent Units; Weighted-Average Method [LO 6-1, 6-2, 6-4]

Washington State Fisheries, Inc., processes salmon for various distributors. Two departments, processing and packaging, are involved. Data relating to tons of salmon sent to the processing department during May follow:

Percent Completed
Tons of Salmon Direct Materials Conversion
Work-in-process inventory, May 1 1,440 80 % 70 %
Work-in-Process inventory, May 31 2,120 50 % 30 %
Started processing during May 7,950

Required:

1. Calculate the number of tons completed and transferred out during the month.

2. Calculate the number of equivalent units for both direct materials and conversion for the month of May, assuming that the company uses the weighted-average method.

3. How would your answer in requirement 2 change if the percentage of completion in ending inventory were as follows: direct materials 30%, conversion 40%?
Answer #1

Solution 1:

Tons completed and transferred out during the month = Work-in-process inventory, May 1 + Started processing during May - Work-in-process inventory, May 31

= 1440 + 7950 - 2120 = 7270 tons

Solution 2:

Equivalent units of production (EUP) - Weighted Average method
Units %Material EUP-Material % Conversion EUP- Conversion
Units transferred out 7270 100% 7270 100% 7270
Units of Ending work in process 2120 50% 1060 30% 636
Total Units 9390 8330 7906

Solution 3:

Equivalent units of production (EUP) - Weighted Average method
Units %Material EUP-Material % Conversion EUP- Conversion
Units transferred out 7270 100% 7270 100% 7270
Units of Ending work in process 2120 30% 636 40% 848
Total Units 9390 7906 8118